x 2 -x-1 = 0 and x>0 then what is the value of x 2 +(1/x 2) =? a)3 b) (1+5+2*5 1/2 ) / 2(1+5 1/2 +2) c) 3+5 1/2 d) 6

x2-x-1 = 0 and x>0 then what is the value of x2+(1/x2) =?
a)3
b) (1+5+2*51/2) / 2(1+51/2+2)
c) 3+51/2
d) 6
 

Grade:9

1 Answers

Hasan Naqvi
8 years ago
x2- x – 1 =0
dividing by x,
x – 1 – 1/x =0
x – 1/x = 1
Squaring both sides
x2 + 1/x2 - 2 = 1
x2 + 1/x2 = 3